A 0.5998 sample of a new compound with a molar mass of 93.12 g/mol was analyzed and found to have 0.2322g of carbon,some hydrogen,and 0.3091g of oxygen. What is the formula for the compund?
Answer #1

% C = (0.2322 / 0.5998) x 100 = 38.7

% O = (0.3091 / 0.5998) x 100 = 51.5

% H = 100 - 51.5 - 38.7 = 9.8

% element atomic ratio simplest ratio 3.225/3.2181.0 C 38.7 38.7/12 3.225 -9.8 98/3.218-3.0 H-9.8 3.218/3.218 1.0 O-51.5 51.5 16 3.218 empirical formulaCH30 empirical formula mass 31 g/mol molar mass-93.12 nmolar mass / empirical formula mass 93.12 31-3.00 molecular formula(CH30)3C3H903